Mr. and Mrs. Jackson invested some money at 6% simple interest and some money at 10% simple interest. In the second investment they put $1000.00 more than they put in the first. If the income from both investments for 1 year was $340.00, then how much did they invest at each rate?
Answer by mananth(16949)

6%--------x
10%---------y
y=x+1000
y-x=1000...........1
6%x+10%y=340
multiply by 100
6x+10y=34000.....2

-1 x + 1 y = 1000 .............1
6 x + 10 y = 34000 .............2
Eliminate y
multiply (1)by -10
Multiply (2) by 1
10 x -10 y = -10000
6 x 10 y = 34000
Add the two equations
16 x = 24000
/ 16
x = 1500
plug value of x in (1)
-1 x + 1 y = 1000
-1500 + 1 y = 1000
1 y = 1000 + 1500
1 y = 2500
y = 2500